To enter a very private garden you need to go through four doors. At each door you must pay an entry fee. If you pay $x at a given door, then you must pay $(2x + 1) at the next door. If it costs a total of $86 to get through the four doors, how much did you pay to get through the first door?

Answer by stanbon(75887)

To enter a very private garden you need to go through four doors. At each door you must pay an entry fee. If you pay $x at a given door, then you must pay $(2x + 1) at the next door. If it costs a total of $86 to get through the four doors, how much did you pay to get through the first door?
---------------
1st: x
2nd: 2x+1
3rd: 2(2x+1)+1 = 4x+3
4th: 2(4x+3)+1 = 8x+7
----------------------------
Solve: x + 2x+1 + 4x+3 + 8x+7 = 86
15x + 11 = 86
15x = 75
x = $5
=============
Cheers,
Stan H.
==============